Assume that you don't start saving for retirement until age 40. You now have 15 years to accumulate 2,500,000.

How much would you have to save each year to accumulate this total? Assume you invest in the stock market generating a rate of return of 10%
Answer #1

Assuming annual saving are invested at end of period

Annual savings *FVA10%,15 = Future value

X * 31.77248 = 2,500,000

X = 2,500,000 / 31.77248

X = 78,684.45

Annual savings = 78,684.45

**Note :If savings are invested at beginning of period then Future value =annual savings *FVAD10%,15

                                                                                                    2,500,000 = Y * 34.94973

                                                             Annual saving = 2500000/34.94973

                                                                                     = 71531.31

Find future value annuity factor from table at 10% for 15 period or using financial calculator where i =10% n= 15 and PMT =1